OpenStudy (anonymous):

any calculus guys out there? how do you find the first derivative of z^2+1/3z

OpenStudy (anonymous):

derivative of z^2 is 2x. derivative of 1/z is -1/z^2 You should be able to do the rest.

OpenStudy (anonymous):

\[\huge \frac{\partial (z^2+\frac{1}{3}z)}{\partial z} = 2z+\frac{1}{3}\]

OpenStudy (diyadiya):

d/dz(z^2) + d/dz(1/3z) 2z-1/3z^2

OpenStudy (anonymous):

f' = 2z + (1/3) * - z^-2 = 2z - 1/3z^2

OpenStudy (anonymous):

\[\huge \frac{\partial (z^2+\frac{1}{3}z^{-1})}{\partial z} = 2z-\frac{1}{3}z^2\]

OpenStudy (anonymous):

For some number n, \[\huge \frac{\partial (z^n)}{\partial z} = nz^{n-1}\]

OpenStudy (anonymous):

i guess the question is, is this a local toy store sells whistles in packages of 3,6,or 12. if you need 30 whistles for a party, in how many diffrent combinations can you buy them? \[z^2+\frac{1}{3}z\] or \[z^2+\frac{1}{3z}\]

OpenStudy (anonymous):

cyter, the symbols you are using should be reserved for partial derivatives. A plain script d is the application here.

OpenStudy (anonymous):

in the back of my book it says that the answer is (z^2-1)/3z^2. i just don't understand how to get that answer but i know that i am supposed to use the power rule

OpenStudy (anonymous):

that is because ( i am just guessing) that your problem is \[f(z)=\frac{z^2+1}{3z}\]

OpenStudy (anonymous):

@satellite73 - what syntax tool do you use to write equations in that format? Much cleaner to see than flat text writing.

OpenStudy (anonymous):

no one actually answered that question. the answers here were to \[\frac{d}{dz}[z^2+\frac{1}{3}z]\] and \[\frac{d}{dz}[z^2+\frac{1}{3z}]\]

OpenStudy (anonymous):

@gt latex

OpenStudy (anonymous):

Use the quotient rule

OpenStudy (anonymous):

ok i solved it using the quotient rule, but since im guessing my calc teacher will have me solve a problem like that using only the power rule on the test. so can anyone solve it using the power rule and getting the correct answer that i got in my book, and show me how to do it?

OpenStudy (anonymous):

ok it would really help if we knew exactly what the problem was

OpenStudy (anonymous):

to find the derivative of that function, you must use both the power rule and the quotient rule

OpenStudy (anonymous):

i repeat, it would really be helpful to know what "that function" actually is. maybe we can use nothing but the "power rule"... maybe